OpenStudy (anonymous):

simplify. assume all variables represent non-zero numbers. the answer should not contain negative exponents. (3d^5)(-7d^8)

OpenStudy (calculusxy):

\(\large 3d^5 \times -7d^8\) I like to group this into \(3 \times -7\) and \(d^5 \times d^8\). I am not sure if this approach is correct, however. But if I do do that then this would be \(-21d^{13}\).

OpenStudy (calculusxy):

@Nnesha Would this approach be correct?

Nnesha (nnesha):

yes. when we multiply same bases we should add the exponents and multiply coefficients